OpenStudy (anonymous):

My answer is B Solve 2x + 3 = 7 (2 points) Answer x ≈ -5.19 x ≈ 0.94 x ≈ -0.19 x ≈ 2.81

OpenStudy (anonymous):

x=2 :/

OpenStudy (anonymous):

Minus three from both sides, then divide both sides by 2. Then \(x=2\)

OpenStudy (campbell_st):

is this a log question..?

OpenStudy (anonymous):

The answer is D because it is approximately 2.

OpenStudy (anonymous):

yeah I got x = 2 as well but this is Solving Exponential Equations with Unequal Bases

OpenStudy (campbell_st):

so the question is \[2^{x + 3} = 7\]

OpenStudy (anonymous):

yes

OpenStudy (johnweldon1993):

ahhh...a log question...that makes sense lol...take the logs of both sides (x+2) log2 = log 7 distribute the log2 then simplify

OpenStudy (campbell_st):

take the log of both sides and using the lag laws for powers you get \[(x + 3) \log(2) = \log(7)\] divide both sides by log(2) \[x + 3 = \frac{\log(7)}{\log(2)}\] now subtract 3 from both sides \[x = -3 + \frac{\log(7)}{\log(2)}\] just type it into a calculator for the answer

OpenStudy (johnweldon1993):

sorry for the typo...i meant (x+3) do as @campbell_st has shown :)

OpenStudy (anonymous):

-0.19264507794

OpenStudy (campbell_st):

thats correct so it looks like the 3rd choice

OpenStudy (anonymous):

thank you xo